Perth Mining Company operates two mines for the purpose of extracting gold and silver. The Saddle Mine costs $15,000/day to operate, and it yields 50 oz of gold and 3000 oz of silver each of x day. The Horseshoe Mine costs $20,000/day to operate, and it yields 75 oz of gold and 1000 oz of silver each of y day. Company management has set a target of at least 650 oz of gold and 18,000 oz of silver. How many days should each mine be operated so that the target can be met at a minimum cost?

(x, y) =   



What is the minimum cost?
$
Answer #1

Please thumbs up

650 3000 n + 1000 y 3 18,000 (60), (0, 18). Saddle Mine & Storseshoe Target minely gold 50 75 šilues. 3000 1ooo 18,000 Cost 13000n + 1000y? 18,000 (0,18) AR 6076-687) 5on +757650 w CS Scannec-with CamScanner (6,0)3000n + 4500g 3500 y = point B sont 754 -650x60 39,000 -3000n + 18,000 = 21,000 = 6 y N = 650 – 750€ 50 B(4,6) cost CAD =a.15